OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points 4,5 and -3,-1 ?

OpenStudy (kamibug):

Point-slope form is: y - y1 = m(x - x1) First use the slope formula to find the slope: -1-5 / -3-5 = -6/-8 = 6/8 = 3/4 m = 3/4 Just plug in the coordinates and the slope to the equation! :) You only need to pick one coordinate since we have the slope. So the equation can be either of the following: y - 5 = 3/4(x - 4) or y + 1 = 3/4(x + 3) Hope this helps! :)